A 333 inch board is to be cut into 3 pieces so that the 2nd piece is 3 times as long as the first piece and the 3rd piece is 5 times as long as the first piece. Find lengths of all 3 pieces..
Answer by checkley77(7072)

Let x be the length of the first piece.
3x is the length of the second piece.
5x is the length of the third piece.
x+3x+5x=333
9x=333
x=333/9
x=37 inches is the length of the first piece.
3*37=111 inches is the length of the second piece.
5*37=185 inches is the length of the third piece.
Proof:
37+111+185=333
333=333